OpenStudy (anonymous):

Find the total area for the regular pyramid. The answer I got was 4+8square root sign 13. I will post a pic

OpenStudy (anonymous):

I don't know what your given are... but I think this will help :) http://math.about.com/od/formulas/ss/surfaceareavol_5.htm

OpenStudy (anonymous):

watch the link :)

OpenStudy (anonymous):

\[\LARGE V=\frac 13 \cdot (2cm)^2\cdot 3cm\] \[\LARGE V=\frac 13 \cdot 4cm^2\cdot 3cm\] \[\LARGE V=\frac 13 \cdot 12cm^3\] \[\LARGE V=4cm^3\]

OpenStudy (anonymous):

I am looking for the total Area not the surface area. I am so lost as to how to get the answer. They are asking for the answer in square root form, and I cannot get it into that form it being wrong.

OpenStudy (anonymous):

auff.. sorry I solved volume . My bad (didn't read question properly.) .. but I think you have the area in the link too... in your case you have to work something... for finding the Area. find some sides, by using pythagoras theorem. and you have to find side "S" (as it is shown in the link) once you have that... everything is over...